OpenStudy (anonymous):

what is the mising value of the answer of the completing the square: comleting square x^2+10x-7=0 answer: x=-5(=/+)4 ( ) missing value on the end is squareroot

OpenStudy (anonymous):

wow another repeat

OpenStudy (anonymous):

yes because the last answer was wrong.. cant figure out the missing value

OpenStudy (anonymous):

here we go again i write it all out \[x^2+10x-7=0\] \[x^2+10x=7\] \[(x+5)^2=7+25=32\] \[x+5=\pm\sqrt{32}=\pm4\sqrt{2}\] \[x=-5\pm4\sqrt{2}\]

OpenStudy (anonymous):

thank you much. i didnt get your answer on the last one? sorry about that

OpenStudy (anonymous):

the last step because \[\sqrt{32}=\sqrt{16\times 2}=\sqrt{16}\times \sqrt{2}=4\sqrt{2}\]

OpenStudy (anonymous):

yw